OpenStudy (lexipoo1998):

Please help. Medal? What is the sum of a 6–term geometric series if the first term is 6 and the last term is 966,306?

OpenStudy (lexipoo1998):

@ganeshie8 @jdoe0001 @cwrw238 can any of you help?

OpenStudy (lexipoo1998):

@nincompoop @nphuongsun93 please help me with this somebody.

OpenStudy (cwrw238):

first find the common ration awe know the first and last terms so ar^5 / a = r^5 = 966,306 / 6 = 161051 now you need the 5th root of 161051 can u use your calculator to find this?

OpenStudy (lexipoo1998):

how would i plug that into my caculator?

OpenStudy (cwrw238):

161,051 ^ 0.2

OpenStudy (cwrw238):

got it?

OpenStudy (lexipoo1998):

it came out to 11

OpenStudy (cwrw238):

then use the formula for sum of n terms Sn = a1 * r^n - 1 -------- r - 1 where a1 = first term and r = common ration

OpenStudy (cwrw238):

11 is correct

OpenStudy (cwrw238):

so sum 6 terem = 6 * (11^6 - 1) -------- 11 - 1

OpenStudy (lexipoo1998):

10629360?

OpenStudy (cwrw238):

no - I think you've missed out the last division by 10

OpenStudy (cwrw238):

right?

OpenStudy (lexipoo1998):

oh oops so it would be 1062936?

OpenStudy (cwrw238):

that's close - I made a slight mistake in the last formula 6 * (11^6 - 1) ------------ 11 - 1 isn't correct - the parentheses should not be there correct is 6 * 11^6 - 1 --------- 11 - 1 That works out to 10629366 - 1 ------------ = 1062936.5 10

OpenStudy (lexipoo1998):

Ok thank you!